# Hallucination, reliability, and the role of generative AI in science

### v1

@@ 4 AlphaFold and the neutralization of hallucination
This admission makes clear that hallucination is not a marginal failure mode, but a central epistemic challenge for scientific AI. So how does the AlphaFold 3 team avoid the problem of corrosive hallucination? The answer lies in two main strategies: (i) the use of theoretical knowledge to guide training, and (ii) the use of confidence-based error screening to guide the interpretation of model output.
@@ 4.1 Theory-guided training
Unlike large language models, [^4] Some generative protein folding models, such as Meta’s ESM protein model, are described by their authors as “language models”, despite being trained on biochemical data rather than natural language. When I use the term “language model”, I am referring to models trained on natural language. which learn statistical structure from vast, heterogeneous, and poorly organized datasets, AlphaFold 3 is trained on highly curated data by means of procedures that encode well-established physical and biochemical constraints. In particular, these constraints are encoded in carefully designed violation loss functions in the training phase. Candidate outputs are penalized if they exhibit steric clashes (where atoms are closer than the van der Waals radius allows), implausible bond lengths, or physically unrealistic torsional angles. These penalties shape the diffusion model’s learned score function by rewarding outputs that adhere to molecular physics, thereby suppressing structurally incoherent predictions.
The epistemic force of AlphaFold’s constraint-based design stems not only from the content of the physical laws it encodes, but from the fact that the evidential basis for those laws is largely independent of the training distribution. Confirmation is strengthened when distinct bodies of empirical knowledge, each grounded in a different measurement techniques, converge on a common target [Sober, 1989]. The structural regularities distilled into the PDB and the theoretical constraints operationalized in the loss function arise from separate physical processes and measurement paradigms. Their convergence in AlphaFold’s architecture transforms an inductive generalization into a theoretically disciplined scientific inference.
@@
This analysis also calls for a reconsideration of a well-known critique of deep learning. Pearl and Mackenzie [2018] and Marcus and Davis [2019] argue that deep learning models cannot acquire causal knowledge. AlphaFold’s design challenges that view. Theoretical constraints, grounded in physics and chemistry, actively shape the model’s optimization process. Moreover, its training data—the Protein Data Bank—is not a random sample of molecular configurations. It is a theoretically curated archive of structures inferred through techniques like X-ray crystallography, cryo-EM, and NMR spectroscopy. At least some of this causal knowledge is embedded in AlphaFold’s learned latent space. So, although AlphaFold may not represent mechanisms explicitly, its training is disciplined by physical law, and its outputs reflect constraints grounded in causal structure. Insofar as the AlphaFold team has built a reliable workflow, that reliability is not a mysterious byproduct of scale but a consequence of carefully engineered design that is guided by knowledge of causal mechanisms at the molecular scale.
@@ 4.2 Confidence-based error screening
The central tool here is the Predicted Local Distance Difference Test (pLDDT). Rather than measuring proximity to training examples, pLDDT estimates the local reliability of a predicted structure based on internal consistency cues. Specifically, AlphaFold generates multiple structure predictions through a stochastic sampling procedure, and pLDDT scores reflect the degree of local agreement among these samples. Where predictions converge tightly, the model assigns high confidence; where they diverge—often due to physical indeterminacy or lack of constraint—it flags the output as unreliable. The underlying idea is simple but powerful: hallucinations are not uniform across stochastic samples. By generating multiple outputs with different random seeds, AlphaFold can identify regions of disagreement and treat them as signals of uncertainty. Idiosyncratic errors tend to cancel out in the aggregate, allowing the model to screen for instability without requiring access to ground truth.666pLDDT scores are produced by a head in AlphaFold’s architecture that is trained to predict the expected deviation between predicted and true interatomic distances for each residue. During training, this head is supervised using experimentally validated structures, allowing the model to calibrate its internal confidence estimates. At inference time, however, the score is computed purely from the model’s own internal representations; no comparison to ground truth is made.
@@
Crucially, AlphaFold achieves this level of reliability despite the opacity of its internal representations. Its trustworthiness stems not from interpretability, but from the way it embeds domain knowledge, performs confidence-based screening, and renders that information in a form that enables downstream decision making about which experiments are most likely to succeed. This is the core insight of computational reliabilism [Durán and Formanek, 2018; Duran, 2023]: models can support reliable inference even when they are not transparent, so long as they are embedded in a well-designed error-screening workflow.
@@ 5 From molecules to meteorology
If AlphaFold demonstrates how generative AI can support inference in molecular biology, GenCast shows how the same epistemic principles extend to large-scale dynamical systems. Unlike protein folding, where the goal is to infer a stable conformational structure, meteorological forecasting targets a dynamic and chaotic system in which small errors in initial conditions can rapidly amplify [Lorenz, 1995], and which lacks a clearly defined end state. Traditional numerical weather prediction (NWP) systems, such as those used by the European Centre for Medium-Range Weather Forecasts (ECMWF), generate predictions by numerically solving fluid dynamics equations to produce ensemble forecasts. These methods are physically grounded and benefit from interpretable parameters, but they are also computationally expensive. In settings where the goal is to predict intrinsically low-frequency, high-impact events—such as floods or wildfires—which lie in the tails of the probability distribution, an extremely large ensemble is required to achieve adequate coverage, and computational costs grow rapidly with event rarity.
GenCast offers a (relatively) computationally efficient alternative. It is a new, diffusion-based generative model trained on ERA5, a reanalysis dataset [^7] In meteorology, a reanalysis is a dataset created by filling gaps in historical observations using numerical models constrained by physical laws. The result is a physically coherent estimate of past atmospheric states, widely used for climate research, model training, and long-term forecasting.that integrates observational data with physics-based simulations. Like AlphaFold, GenCast learns the statistical structure of valid trajectories and generates plausible forecasts via a generative process. Though it lacks an explicit representation of fluid dynamics, GenCast now matches or exceeds the performance of ECMWF’s gold-standard system on several standard forecasting metrics (and is, moreover, the first machine-learning weather prediction model to rival the accuracy of traditional simulation-based approaches) [Price et al., 2024].
@@
GenCast’s success, like AlphaFold’s, depends on its being embedded in a workflow that draws on independent scientific theory. It is trained not on raw observations but on ERA5: reconstructions that obey conservation laws and reflect established meteorological principles. This mirrors AlphaFold’s reliance on the Protein Data Bank, which functions as a theoretically curated archive of molecular structure. Second, like AlphaFold, GenCast incorporates physics-informed loss functions that penalize predictions violating basic physical constraints [Kashinath et al., 2021]. Earlier machine learning models could generate forecasts that appeared locally plausible but violated global coherence. For example, they might predict a negative humidity value or a physically unrealistic temperature gradient [Watt-Meyer et al., 2021]. GenCast avoids such failures by incorporating domain-specific constraints during training and by relying on architectures that tend to preserve these constraints at inference time. These practices reduce the risk that model outputs will be epistemically disruptive, in the sense defined in Section 3.
@@
Like AlphaFold, GenCast addresses hallucination through confidence-based error screening, though by different means. Instead of assigning explicit confidence scores to individual predictions, GenCast introduces stochastic variation at inference time, generating an ensemble of plausible forecasts from different random seeds. Given the chaotic nature of weather systems, each trajectory varies in local details, but the ensemble as a whole preserves coherent global structure. The epistemic value of this approach lies in how this variability reveals where inference is likely to be unreliable. Because hallucinations differ across stochastic runs, their dispersion serves as a signal of epistemic instability. Unstable predictions appear as outliers, while robust features emerge as recurring patterns. In this way, GenCast provides a confidence signal. Although it is not a separately computed output, as it is in AlphaFold, it is a reliable statistical pattern that expert scientists can leverage.
@@
One way to assess how well GenCast’s internal uncertainty estimates align with forecasting performance is through the spread–skill ratio, which compares the ensemble’s internal variance (the spread) with its actual forecast error (the skill). [^8] The spread–skill ratio (SSR) compares an ensemble’s internal variance (spread) to its forecast error (root-mean-square error, RMSE). It is given by: $\text{SSR}=\frac{\text{Spread}}{\text{RMSE}}=\frac{\sqrt{\frac{1}{N}\sum_{i=1}^{N}(f_{i}-\bar{f})^{2}}}{\sqrt{\frac{1}{M}\sum_{j=1}^{M}(\bar{f}_{j}-o_{j})^{2}}}$ where $f_{i}$ is the forecast from ensemble member $i$, $\bar{f}$ is the ensemble mean forecast, $o_{j}$ is the observation at verification case $j$, $N$ is the number of ensemble members, and $M$ is the number of forecast–observation pairs. An SSR near 1 indicates well-calibrated uncertainty; values significantly greater or less than 1 suggest overdispersion or underdispersion, respectively [Fortin et al., 2014]. Variations on this definition exist, and there is ongoing debate about which formulation is most appropriate in different settings.A spread–skill ratio near 1 indicates that the model’s uncertainty estimates are well-matched to its performance—neither overconfident nor needlessly conservative. In the GenCast evaluation, this ratio remained close to 1 across a range of forecast horizons (i.e., the time intervals into the future for which predictions are made), confirming that the ensemble’s internal dispersion reliably mirrors the inherent empirical uncertainty in chaotic systems.
@@
GenCast’s success shows how hallucinations can be made epistemically tractable. Rather than being left as corrosive, errors are converted into expected, bounded deviations whose impact can be absorbed by the broader workflow. This undermines the suspicion that opaque generative models leave us with no alternative but what Duede memorably called brute inductivism. On that view, the lack of interpretability precludes the possibility of theory-guided inference, and users are left to trust outputs solely on the basis of observed empirical correlations. But as the GenCast case makes clear, we have more to rely on here than the naked predictive track record. By combining theory-informed training, ensemble-based uncertainty modeling, and explicit error detection practices, GenCast enables a mode of inference that is neither brute nor blind. It underscores the same lesson we saw in AlphaFold: generative models become reliable not by virtue of their internal transparency, but by being embedded in theory-rich, uncertainty-aware practices that help scientists anticipate and manage error.
@@ 6 Discovery and justification
This objection is inspired by Duede’s 2023 argument that concerns about AI reliability often reflect a misunderstanding of its scientific role. Duede claims that AI is fundamentally a tool for discovery, not for justification. On this view, my concern about hallucination wrongly presupposes that AI is in the business of delivering justification for model outputs. Duede might argue that AI merely offers heuristic guidance: it narrows the space of relevant hypotheses, but those hypotheses only become candidates for belief once they have been subjected to empirical test. From a thoroughly empiricist standpoint, the strategies scientists use to mitigate hallucination appear secondary—or even unnecessary.
This response echoes Karl Popper’s 1959 appeal to the influential distinction between the context of discovery and the context of justification. Popper famously held that the epistemology of science should focus solely on justification through rational reconstruction, since discovery is largely governed by factors like intuition and creativity, which are, according to Popper, beyond rational control.
Yet the historical turn in the philosophy of science has long challenged the sharp division between discovery and justification. Scientific heuristics, whether human or AI-driven, are not arbitrary guesses; they are assessed by their empirical traction. This is especially evident in the development and funding of generative AI systems like AlphaFold. The model was not funded merely for its capacity to generate intriguing hypotheses, but because its developers demonstrated—before large-scale empirical validation—that it could reliably predict biologically plausible protein structures. Its ability to infer accurate 3D conformations from amino acid sequences had clear implications for understanding biological function and disease. This predictive success led to rapid adoption across the life sciences, where researchers now use its outputs to guide experiment and hypothesis formation. AlphaFold is not treated as a tool for blind exploration, but as a theory-informed model capable of supporting novel inferences. When scientists take one of its outputs to be approximately true, their belief enjoys a kind of partial justification.
@@
More generally, the justification for placing credence in generative AI systems is embedded throughout their scientific workflows, and not merely appended afterward.

### v2

@@ 4 AlphaFold and the neutralization of hallucination
This admission makes clear that hallucination is not a marginal failure mode, but a central epistemic challenge for scientific AI. So how does the AlphaFold 3 team mitigate hallucination? The answer lies in two main strategies: (i) the use of theoretical knowledge to guide training, and (ii) the use of confidence-based error screening to guide the interpretation of model output.
@@ 4.1 Theory-guided training
Unlike large language models, [^6] Some generative protein folding models, such as Meta’s ESM protein model, are described by their authors as “language models”, despite being trained on biochemical data rather than natural language. When I use the term “language model”, I am referring to models trained on natural language. which learn statistical structure from vast, heterogeneous, and poorly organized datasets, AlphaFold 3 is trained on the Protein Data Bank (PDB), a highly curated repository of experimentally validated molecular structures. Moreover, the training procedures encode well-established physical and biochemical constraints through carefully designed violation loss functions. Candidate outputs are penalized if they exhibit steric clashes, implausible bond lengths, or physically unrealistic torsional angles—the rotational angles around chemical bonds that determine backbone geometry.
The necessity of these constraints is demonstrated empirically. The AlphaFold team reports that without violation loss terms, “the network is observed to frequently violate the chain constraint during the application of the structure module” [Jumper et al., 2021]. That is, the model produces structures with impossible bond geometries—steric clashes where atoms are represented closer together than the van der Waals radius permits, peptide bonds at wrong angles, and so on. The violation loss actively suppresses such physically incoherent predictions by shaping the diffusion model’s learned score function to reward outputs that adhere to molecular physics. Even with these penalties in place, AlphaFold’s raw outputs still require a final refinement step using molecular dynamics simulations (Amber force field) to perfectly enforce physical constraints. This post-processing step indicates just how difficult it is to satisfy these constraints through neural network training alone.
The epistemic force of AlphaFold’s constraint-based design stems not only from the content of the physical laws it encodes, but from the fact that the evidential basis for those laws is largely independent of the training distribution. Confirmation is strengthened when distinct bodies of empirical knowledge, each grounded in a different measurement techniques, converge on a common target [Sober, 1989; Schupbach, 2018]. The structural regularities distilled into the PDB and the theoretical constraints operationalized in the loss function arise from separate physical processes and measurement paradigms. Their convergence in AlphaFold’s architecture transforms an inductive generalization into a theoretically disciplined scientific inference.
@@
This analysis also speaks to concerns about whether deep learning models can acquire causal knowledge. Some critics worry that DNNs merely find statistical patterns without learning the causal structure that underlies them [Pearl, 2018; Marcus and Davis, 2019]. AlphaFold’s design suggests a more nuanced picture. While the model may not represent causal mechanisms in a form that supports arbitrary counterfactual reasoning, its training is nevertheless disciplined by causal knowledge. Theoretical constraints grounded in physics and chemistry actively shape the optimization process. Moreover, the Protein Data Bank is not a random sample of molecular configurations but a theoretically curated archive of structures inferred through techniques like X-ray crystallography, cryo-EM, and NMR spectroscopy. These techniques themselves depend on causal models of how electromagnetic radiation interacts with molecular structure. So although AlphaFold may not contain an explicit, manipulable causal model à la Pearl, its learned representations implicitly encode causal constraints from molecular physics.
@@ 4.2 Confidence-based error screening
The central tool here is the Predicted Local Distance Difference Test (pLDDT). Rather than measuring proximity to training examples, pLDDT estimates the local reliability of a predicted structure based on internal consistency cues. Specifically, AlphaFold generates multiple structure predictions through a stochastic sampling procedure, and pLDDT scores reflect the degree of local agreement among these samples. Where predictions converge tightly, the model assigns high confidence; where they diverge—often due to physical indeterminacy or lack of constraint—it flags the output as unreliable. The underlying idea is simple but powerful: hallucinations are not uniform across stochastic samples. By generating multiple outputs with different random seeds, AlphaFold can identify regions of disagreement and treat them as signals of uncertainty. Idiosyncratic errors tend to cancel out in the aggregate, allowing the model to screen for instability without requiring access to ground truth.888pLDDT scores are produced by a head in AlphaFold’s architecture that is trained to predict the expected deviation between predicted and true interatomic distances for each residue. During training, this head is supervised using experimentally validated structures, allowing the model to calibrate its internal confidence estimates. At inference time, however, the score is computed purely from the model’s own internal representations; no comparison to ground truth is made.
@@
Crucially, AlphaFold achieves this level of reliability despite the opacity of its internal representations. Its trustworthiness does not depend on understanding what individual parameters represent, but rather on how the model is embedded in a carefully designed workflow. Theory-guided training shapes the model’s outputs through physical constraints, while confidence-based screening provides systematic signals about when those outputs can support reliable inference. This is the core insight of computational reliabilism [Durán and Formanek, 2018; Duran, 2023]: models can support reliable inference even when they are not transparent, so long as they are embedded in well-designed error-screening workflows.
@@
4.3 Scope and Limitations
The strategies demonstrated here succeed because structural biology is a theoretically mature domain. The Protein Data Bank encodes decades of experimental work, yielding a training corpus of over 200,000 structures. The violation loss functions operationalize physical laws that have been validated across countless experiments: van der Waals radii, bond angle constraints, torsional potentials. Confidence scores can be calibrated against ground truth because experimental validation techniques like X-ray crystallography, cryo-EM, and NMR spectroscopy are well-established and widely available.
In domains where such resources are absent, comparable reliability may be unattainable. Theory-guided training requires robust theory to guide it. Confidence-based screening requires extensive validation data to calibrate against. AlphaFold’s success demonstrates what is possible when generative AI is embedded in a mature scientific framework, but it does not show that similar workflows can be constructed anywhere.
@@ 5 From molecules to meteorology
If AlphaFold demonstrates how generative AI can support inference in molecular biology, GenCast shows how the same epistemic principles extend to large-scale dynamical systems. Unlike protein folding, which targets a stable conformational structure, meteorological forecasting concerns a chaotic, evolving system in which small errors in initial conditions can rapidly amplify [Lorenz, 1995]. It also lacks a clearly defined end state. Traditional numerical weather prediction (NWP) systems, such as those developed by the European Centre for Medium-Range Weather Forecasts (ECMWF), generate forecasts by numerically solving fluid dynamics equations. These methods are physically grounded and benefit from interpretable parameters, but they are computationally intensive. The challenge is especially acute for low-frequency, high-impact events—such as floods or wildfires—which lie in the tails of the distribution. Capturing them reliably requires extremely large ensembles, and computational costs rise steeply with event rarity.
GenCast offers a relatively computationally efficient alternative. It is a new, diffusion-based generative model trained on ERA5, a reanalysis dataset [^9] In meteorology, a reanalysis is a dataset created by assimilating diverse historical observations into numerical weather prediction models, producing a spatially and temporally coherent reconstruction of past atmospheric states. Although physically constrained, reanalysis outputs are model-dependent and may reflect biases or limitations of the underlying data assimilation systems [McGovern et al., 2024]. produced by the European Centre for Medium-Range Weather Forecasts (ECMWF). ERA5 combines physics-based simulations with extensive observational data, including satellite, radar, and ground measurements, and uses advanced data assimilation techniques. The result is not a record of direct observations. Rather, it is a model-informed reconstruction that aims to balance empirical accuracy with physical coherence.
Like AlphaFold, GenCast learns the statistical structure of valid trajectories and generates plausible forecasts via a generative process. Though it lacks an explicit representation of fluid dynamics, it matches or exceeds the performance of traditional simulation-based systems on several standard forecasting metrics [Price et al., 2024]. Earlier models such as Pangu Weather [Bi et al., 2023] achieved comparable accuracy, but GenCast was the first to combine this with a systematic probabilistic evaluation framework. This probabilistic framework enables systematic assessment of when and where the model’s predictions are reliable.
@@
As with AlphaFold, GenCast’s reliability stems from its integration into a theory-informed workflow. Both systems are trained not on raw, unstructured data but on theoretically curated datasets: AlphaFold on the Protein Data Bank, which encodes experimentally validated molecular structures, and GenCast on ERA5, which integrates physical models with observational data to reconstruct coherent atmospheric states. In each case, domain knowledge shapes the training corpus itself, ensuring that the model learns from inputs already disciplined by physical theory. Moreover, like AlphaFold, GenCast incorporates physically grounded loss functions that penalize violations of general physical laws such as conservation of mass, momentum, and energy [Kashinath et al., 2021]. Earlier machine learning models could generate forecasts that appeared locally plausible but violated global coherence. For example, they might predict a negative humidity value or a physically unrealistic temperature gradient [Watt-Meyer et al., 2021]. GenCast avoids such failures by incorporating these physical constraints during training and by relying on architectures that tend to preserve them at inference time. These practices reduce the risk that model outputs will be epistemically disruptive, in the sense defined in Section 3.
@@
These theoretical constraints help ensure that GenCast’s outputs are physically plausible—but plausibility alone does not guarantee reliability. As with AlphaFold, GenCast also implements a strategy for detecting and managing residual errors. In place of AlphaFold’s explicit confidence scores, GenCast addresses hallucination through a form of ensemble-based uncertainty estimation. Rather than assigning confidence values to individual predictions, it introduces stochastic variation at inference time, generating an ensemble of plausible forecasts from different random seeds. Given the chaotic nature of weather systems, each trajectory varies in local details, but the ensemble as a whole preserves coherent global structure—reflecting past variability rather than hand-tuned perturbations [Lessig et al., 2023]. The epistemic value of this approach lies in how this variability reveals where inference is likely to be unreliable. Because hallucinations differ across stochastic runs, their dispersion serves as a signal of epistemic instability. Unstable predictions appear as outliers, while robust features emerge as recurring patterns. In this way, GenCast provides a confidence signal. Although it is not a separately computed output, as it is in AlphaFold, it is a reliable statistical pattern that expert scientists can leverage.
@@
One way to assess how well GenCast’s internal uncertainty estimates align with forecasting performance is through the spread–skill ratio, which compares the ensemble’s internal variance (the spread) with its actual forecast error (the skill). [^10] The spread–skill ratio (SSR) compares an ensemble’s internal variance (spread) to its forecast error (root-mean-square error, RMSE). It is given by: $\text{SSR}=\frac{\text{Spread}}{\text{RMSE}}=\frac{\sqrt{\frac{1}{N}\sum_{i=1}^{N}(f_{i}-\bar{f})^{2}}}{\sqrt{\frac{1}{M}\sum_{j=1}^{M}(\bar{f}_{j}-o_{j})^{2}}}$ where $f_{i}$ is the forecast from ensemble member $i$, $\bar{f}$ is the ensemble mean forecast, $o_{j}$ is the observation at verification case $j$, NN is the number of ensemble members, and $M$ is the number of forecast–observation pairs. An SSR near 1 indicates well-calibrated uncertainty; values significantly greater or less than 1 suggest overdispersion or underdispersion, respectively [Fortin et al., 2014]. Variations on this definition exist, and there is ongoing debate about which formulation is most appropriate in different settings.A spread–skill ratio near 1 indicates that the model’s uncertainty estimates are well-matched to its performance—neither overconfident nor needlessly conservative. In the GenCast evaluation, this ratio remained close to 1 across a range of forecast horizons (i.e., the time intervals into the future for which predictions are made), confirming that the ensemble’s internal dispersion reliably mirrors the inherent empirical uncertainty in chaotic systems [Price et al., 2024].
5.1 Scope and Limitations
GenCast inherits its reliability from meteorology’s theoretical and observational infrastructure. Atmospheric dynamics obey quantitative physical laws (Navier-Stokes equations, thermodynamic principles, conservation laws) that can be encoded in loss functions. The ERA5 reanalysis dataset integrates decades of satellite, radar, and ground observations with physics-based models, providing a training corpus that is both extensive and theoretically constrained. Forecast verification is rapid: predictions can be checked against actual weather within days or weeks, enabling continuous calibration of ensemble-based uncertainty estimates through metrics like the spread-skill ratio. This combination of quantitative theory, dense observational networks, and fast empirical feedback is not universal. In domains lacking these features—where theories are qualitative, data are sparse, or validation timescales are long—ensemble-based uncertainty estimation may be less reliable and physical constraints harder to operationalize. For example, climate modeling on decadal timescales faces validation challenges that weather forecasting does not, since we cannot quickly verify 30-year projections. Similarly, in fields where physical laws are poorly understood or observational coverage is patchy, the confidence signals that GenCast provides may not emerge reliably. GenCast demonstrates how theory-rich workflows can discipline generative inference under favorable conditions, but these conditions cannot be assumed to hold everywhere.
@@ 5.1 Scope and Limitations
GenCast’s success shows how hallucinations can be made epistemically tractable. Rather than causing scientists to adopt false beliefs, errors are converted into expected, bounded deviations that the broader workflow is designed to absorb. This undermines the suspicion that opaque generative models leave us with no alternative but what Duede memorably called brute inductivism. On that view, the lack of interpretability precludes the possibility of theory-guided inference, and users are left to trust outputs solely on the basis of observed empirical correlations. But as the GenCast case makes clear, we have more to rely on here than the naked predictive track record. Ensemble-based uncertainty modeling converts errors into calibrated signals about where predictions are reliable. By combining theory-informed training with explicit error detection practices, GenCast enables a mode of inference that is neither brute nor blind. It underscores the same lesson we saw in AlphaFold: generative models become reliable not by virtue of their internal transparency, but by being embedded in theory-rich, uncertainty-aware practices that help scientists anticipate and manage error.
@@ 6 Discovery and justification
This objection is inspired by Duede’s [2023] argument that concerns about AI reliability often reflect a misunderstanding of its scientific role. Duede claims that AI is fundamentally a tool for discovery, not for justification. On this view, my concern about hallucination wrongly presupposes that AI is in the business of delivering justification for model outputs. Duede might argue that AI merely offers heuristic guidance: it narrows the space of relevant hypotheses, but those hypotheses only become candidates for belief once they have been subjected to empirical test. From a thoroughly empiricist standpoint, the strategies scientists use to mitigate hallucination appear secondary—or even unnecessary.
This response echoes Karl Popper’s [1959] influential distinction between the context of discovery and the context of justification. Popper famously argued that the epistemology of science should concern itself solely with justification through rational reconstruction, since the processes of discovery are guided by intuition, creativity, and other factors beyond rational control.
Yet the historical turn in the philosophy of science has cast doubt on the sharpness of this division. Popper’s distinction, while conceptually useful, is ultimately artificial. In practice, discovery and justification are often intertwined. Scientific heuristics are not arbitrary guesses; they are evaluated by their empirical traction and shaped by theoretical expectations. This is especially evident in the development and funding of generative AI systems like AlphaFold. The model was not funded merely for its capacity to generate intriguing hypotheses, but because, prior to large-scale empirical validation efforts, its developers demonstrated that it could reliably predict biologically plausible protein structures. Its ability to infer accurate 3D conformations from amino acid sequences had clear implications for understanding biological function and disease. This predictive success led to rapid adoption across the life sciences, where researchers now use its outputs to guide experimentation and hypothesis formation. AlphaFold is not treated as a tool for blind exploration, but as a theory-informed model capable of supporting novel inferences. When scientists take one of its outputs to be approximately true, their belief enjoys a measure of epistemic justification.
@@
More generally, this suggests a modest lesson: generative models are most effective where background knowledge is sufficiently extensive to constrain their outputs and structure their use. In domains like protein folding or meteorology, theory provides a framework that helps identify and account for error. Where such knowledge is lacking, errors are harder to interpret and more likely to mislead. Generative models can accelerate discovery, but they do so most reliably where prior understanding already runs deep.